OpenStudy (anonymous):

How would I find f prime a for the function f(x)=4x^-2? I have no clue how to set it up.

OpenStudy (amistre64):

still going to be the power rule .... define the power rule for me.

OpenStudy (amistre64):

notifs didnt tell me you replied to the other one

OpenStudy (amistre64):

what have you learned about finding a derivative so far? limit stuff, or are you on to the rules

OpenStudy (anonymous):

I know how to find the derivative, and I THINK i should be using (f(a+h)-f(a))/h but I don't know how to fit the problem in to it. How do I get rid of the ^-2? would it become ^1/4?

OpenStudy (amistre64):

recall that a^(-n) = 1/(a^n) that will help us with it

OpenStudy (amistre64):

\[\frac{(x+h)^{-2}-(x)^{-2}}{h}\] \[\frac{\dfrac{1}{(x+h)^{2}}-\dfrac{1}{x^{2}}}{h}\] then it works our by adding the fractions and working a conjugate most likely

OpenStudy (anonymous):

wouldn't it be 4(x+h)^-2?

OpenStudy (amistre64):

the 4 is a constant, so its really immaterial. if you want to include it feel free

OpenStudy (anonymous):

Ok. Thank you so much!

OpenStudy (anonymous):

I think I can work it out on my own for the mos part now.

OpenStudy (amistre64):

[4 f(x)]' = 4 f'(x) constants pull out

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

YES I got it! Thank you!